Question 2 (2 points) Random samples from two age groups of brides (200 brides under 18 years old and 100 brides showed that 50% of brides in the under 18 group were divorced after 15 years, while 40% of brides n the 20 or older group were divorced after 15 years. The difference between the two proportions is 0.10 with a standard error of 0 is a 95% confidence interval for the difference between the population proportions of brides who are divorced within 15 years (bridies under 18 - brides at least 20)? at least twenty years old) age .0604. What (-0.040,0.160) O none of the above O (-0.123,0.023) (-0.018, 0.218) Save
Answer #1

The confidence Interval for difference in proportions is D \pm Zc*SE

D = 0.1, Zc = 1.96 and SE = 0.0604

The lower Limit = 0.1 - (1.96*0.0604) = -0.018

The upper Limit = 0.1 + (1.96*0.0604) = 0.218

The 4th Option (-0.018,0.218)